### v1

# Tabular foundation model to detect empathy from visual cues

Abstract.
Detecting empathy from video interactions is an emerging area of research. Video datasets, however, are often released as extracted features (i.e., tabular data) rather than raw footage due to privacy and ethical concerns. Prior research on such tabular datasets established tree-based classical machine learning approaches as the best-performing models. Motivated by the recent success of textual foundation models (i.e., large language models), we explore the use of tabular foundation models in empathy detection from tabular visual features. We experiment with two recent tabular foundation models – TabPFN v2 and TabICL – through in-context learning and fine-tuning setups. Our experiments on a public human-robot interaction benchmark demonstrate a significant boost in cross-subject empathy detection accuracy over several strong baselines (accuracy: $0.590\rightarrow 0.730$; AUC: $0.564\rightarrow 0.669$). In addition to performance improvement, we contribute novel insights and an evaluation setup to ensure generalisation on unseen subjects in this public benchmark. As the practice of releasing video features as tabular datasets is likely to persist due to privacy constraints, our findings will be widely applicable to future empathy detection video datasets as well.
empathy, tabular data, foundation model, empathy detection, human-robot interaction, video features
ccs: Computing methodologies Learning paradigmsccs: Applied computing Psychologyccs: Computing methodologies Neural networks
1. Introduction
Figure 1. Overview of the task and motivation. We study the task of detecting empathy of subjects listening to narrative storytelling, where the listener self-reports their empathy level. From the listener’s video, visual features (e.g., eye gaze, facial expressions) are released as a tabular dataset (Mathur et al., 2021). In this context, we propose using tabular foundation models. Our results show that both in-context learning and fine-tuning approaches improve performance, and fine-tuning provides stronger generalisation to unseen subjects.
Empathy, as defined in social psychology (Davis, 1996; Hoffman, 2000), refers to an individual’s capacity to perceive, comprehend and vicariously experience the emotional state of others. This interpersonal skill has practical significance in a variety of interaction contexts, such as human-robot interactions, clinical communication and educational settings. Training intervention to nurture this skill requires measurement [^1] We use the terms ‘measurement’, ‘detection’ and ‘prediction interchangeably. of empathy. Current assessment approaches rely on self-reported questionnaires or labour-intensive manual analysis by external observers (Bas-Sarmiento et al., 2020). Automated empathy detection through machine learning techniques has gained momentum over the past decade (Hasan et al., 2024b). While the majority of this progress has occurred within natural language processing contexts, i.e., detecting empathy in textual content, empathy detection from video data remains relatively under-explored (Hasan et al., 2024b). Towards this direction, this paper measures empathy from video-based behavioural cues.
Several empathy detection benchmarks are proposed in the literature aiming at different aspects of empathy, such as assessing the level of empathy in an interaction (Buechel et al., 2018; Mathur et al., 2021), measuring empathic similarity between two contents (Shen et al., 2023), measuring the changes of someone’s emotional state (Barros et al., 2019), and detecting the direction of empathy (Hosseini and Caragea, 2021). Our work is grounded on the assessment of the level of empathy, being the most common form in the literature (Hasan et al., 2024b). In particular, we detect the level of empathy elicited by storytelling, i.e., measuring how much empathy a listener demonstrates towards others portrayed in a story (Mathur et al., 2021; Keen, 2006) (Figure 1).
Being an emerging area of research, the amount of public video datasets is very limited (Hasan et al., 2024b). Datasets of raw video footage featuring people’s activity are difficult to make public due to privacy and ethical concerns. Consequently, the majority of these few existing public datasets, such as human-robot interactions (Mathur et al., 2021; Spitale et al., 2022), are released as extracted feature vectors (i.e., tabular data) instead of raw video footage. Since privacy and ethical concerns are likely to persist, one of the key challenges in developing an empathy detection system from video is machine learning from tabular datasets.
Traditional approaches to tabular data have long been dominated by tree-based methods like gradient-boosted decision trees (Hollmann et al., 2023, 2025). While deep learning has revolutionised learning from texts and images, it has struggled with tabular data due to the heterogeneity between datasets and the heterogeneous nature of the raw data itself. Foundation models like large language models (LLMs) and large vision models (LVMs) have been widely explored in downstream tasks on texts and videos. Similarly, LLMs have also been explored with tabular data (Hegselmann et al., 2023; Fang et al., 2024), but these approaches convert table rows into some textual format, such as sentences. As LLMs do not natively support numerical tabular data, a new paradigm of the foundation model has emerged: the tabular foundation model. The first one of its kind, a prior data-fitted network (PFN) (Müller et al., 2022) trained on lots of synthetic tabular data, is TabPFN (Hollmann et al., 2023). Following the success of TabPFN, TabICL (tabular in-context learning) (Qu et al., 2025) is another tabular foundation model, introduced in 2025.
Among the few video-based public empathy detection benchmarks, Mathur et al. (2021) features empathy elicited by a robot storyteller. In this dataset, human subjects attend to a robot telling different stories that may elicit empathy in the human subjects. Although there are multiple sessions per subject, prior work has failed to consider scenarios where the same subject’s data may appear in both training and test sets. Given that empathy is subjective, the same subjects appearing in both training and test splits do not ensure true generalisation performance on unseen subjects. To address this, we design experiments and report results in a new evaluation scheme on this benchmark dataset where it is ensured that the same subject’s data is not shared between the training and test sets. Our proposed fine-tuning of the tabular foundation model achieves impressive results on this cross-subject generalisation.
Previous work on video-based elicited empathy detection demonstrated classical machine learning models as state-of-the-art approaches. In contrast, we explore the potential of tabular foundation models and outperform existing methods by a significant margin (Section 4.3 and Section 4.4). Our key contributions are threefold:
  1. (1)
    We design and evaluate in-context learning (ICL) with two tabular foundation models (TabPFN and TabICL) for empathy detection from visual cues.
  2. (2)
    Our proposed fine-tuning of the popular TabPFN foundation model results in further performance improvement, especially in detecting empathy of unseen subjects.
  3. (3)
    We reveal limitations in current evaluation protocols in a public empathy detection benchmark and propose a fairer and unbiased evaluation to ensure robust empathy detection.
2. Related work
2.1. Empathy detection
There are several empathy detection studies in text domains, such as written essays (Buechel et al., 2018; Hasan et al., 2023), social media conversations (Khanpour et al., 2017) and email correspondence (Sedefoglu et al., 2024). In detecting empathy from videos, Mathur et al. (2021) released a human-robot interaction (HRI) dataset, aiming to measure empathy elicited by a robot storyteller. They evaluated 10 different machine learning approaches, including eight classical models (e.g., AdaBoost, bagging, decision trees, support vector machine, logistic regression, random forest and XGBoost) and two deep learning architectures (long short-term memory network and temporal convolutional network). For classical machine learning approaches, features were represented as fixed-length vectors of statistical attributes (mean, median, standard deviation and autocorrelation). For deep learning approaches, the features were maintained as temporal sequences. Through 5-fold stratified cross-validation with 10 repeats, XGBoost appeared as the highest-performing classifier. Contrary to their work, we propose in-context learning and fine-tuning of tabular foundation models in detecting elicited empathy.
Other studies on video-based empathy detection include interaction between human subjects and virtual agents (Kroes et al., 2022; Tavabi et al., 2019), teacher and students (Pan et al., 2022), human subject and avatar (Hervás et al., 2016), counsellor and clients (Zhu et al., 2023), two human subjects (Barros et al., 2019), and human and AI agents (Shen et al., 2024). In contrast to these empathy modelling approaches, we focus on measuring elicited empathy from the public HRI benchmark.
2.2. Application of foundation models
Textual foundation models (i.e., LLMs) have been explored in several stages of empathy detection from textual data. For example, Hasan et al. (2025) used GPT-4 and Llama-3 LLMs and Kong and Moon (2024) used GPT series LLMs in a zero-shot setting to predict empathy in written essays. On the same task, Kong and Moon (2024) and Hasan et al. (2024a) explored few-shot prompt engineering (i.e., in-context learning) with GPT LLMs. Apart from the prediction, LLMs are also explored in data augmentation (Lu et al., 2023; Frick and Steinebach, 2024) and data annotation (Hasan et al., 2025, 2024a).
Motivated by the success of LLMs in empathy detection from text, we explore two recent tabular foundation models released in 2025 – TabPFN v2 (Hollmann et al., 2025) and TabICL (Qu et al., 2025) – for tabular video features. While tabular foundation models’ in-context learning has been demonstrated in numerical datasets such as OpenML test suites (Vanschoren et al., 2014), we explore its capability in dealing with video features. In addition to the in-context learning, we fine-tune TabPFN v2, which further boosts empathy classification performance.
3. Method
Figure 2. Overview of the proposed pipeline for empathy detection using tabular foundation models. Temporal multimodal features (e.g., facial landmarks, eye gaze) are first extracted for each sample $\mathbf{X}_{i}\in\mathbb{R}^{T_{i}\times d}$. These sequences are summarised by computing $n$ statistical attributes (e.g., mean, median) over time, resulting in a fixed-size tabular representation $\mathbf{X}_{i}\in\mathbb{R}^{n\times d}$. This representation is used for binary empathy classification $y_{i}\in\{0,1\}$) via either (top) in-context learning, where a frozen tabular foundation model predicts labels with a single forward pass over the training data, or (bottom) fine-tuning of the foundation model with frozen X and Y encoders. The model predicts whether a subject is empathic or less empathic.
Figure 2 presents an overview of our approach for empathy detection using tabular foundation models. Temporal multimodal features, such as facial landmarks and eye gaze, are first aggregated into fixed-size tabular representations via statistical summarisation. These are then used for binary empathy classification through either in-context learning (ICL) or fine-tuning of a tabular foundation model. Further details of each component are provided in the following subsections.
3.1. Problem formulation
We formulate empathy detection in human-robot interaction videos as a supervised or in-context learning problem. Given a dataset $\mathcal{D}=\{(\mathbf{X}_{i},y_{i})\}_{i=1}^{N}$, consisting of $N$ interaction samples from $M$ subjects, where $\mathbf{X}_{i}\in\mathbb{R}^{T_{i}\times d}$ represents a $d$-dimensional feature vector across $T_{i}$ frames of the $i$-th video sample and $y_{i}\in\{0,1\}$ denotes the corresponding binary empathy label.
The feature vector $\mathbf{X}_{i}$ contains frame-level visual behavioural cues (e.g., eye gaze directions, facial action unit intensities, facial landmarks, and other non-verbal indicators) extracted using toolkits such as OpenFace (Amos et al., 2016). These temporal features can be processed in two ways: (1) time-series modelling or (2) transformed into fixed-length representations by computing statistical attributes (e.g., mean, median, standard deviation and autocorrelation) across the temporal axis. Prior work (Mathur et al., 2021) demonstrated that the latter approach works better for empathy detection in the HRI context.
For $n$ statistical attributes, the input feature representation becomes a fixed-size representation $\mathcal{X}_{i}\in\mathbb{R}^{n\times d}$. To reduce dimensionality, we select $d^{\prime}$ features from the original $d$ features using ANOVA F-values computed between each feature and the target labels. Given this tabular representation, our goal is to learn a function $f:\mathbb{R}^{n\times d^{\prime}}\rightarrow[0,1]$ that maps the input features to a conditional probability distribution over the target variable (empathy level). For a general supervised learning task, it becomes:
(1) $p(y=1|\mathbf{x};\theta)=f_{\theta}(\mathbf{x})$
where $\mathbf{x}\in\mathcal{X}$ represents the input features, and $\theta$ denotes the model parameters optimised during training.
3.2. Traditional supervised learning approaches
In supervised learning, we use labelled training data $\mathcal{D}_{\text{train}}$ to optimise model parameters:
(2) $\theta^{*}=\arg\min_{\theta}\mathcal{L}(\theta,\mathcal{D}_{\text{train}})$
where $\mathcal{L}$ is a loss function (e.g., binary cross-entropy) and $\theta^{*}$ represents the optimised parameters. During inference, predictions are made on a separate test data:
(3) $p(y=1|\mathbf{x};\theta^{*})=f_{\theta^{*}}(\mathbf{x})$
3.3. In-context learning with tabular foundation model
In-context learning explicitly utilises the training data in the inference framework:
(4) $p(y=1|\mathbf{x},\mathcal{D}_{\text{train}};\theta)=f_{\theta}(\mathbf{x},\mathcal{D}_{\text{train}})$
Tabular foundation models such as TabPFN and TabICL support in-context learning. These foundation models are pre-trained on millions of synthetic tabular datasets to predict labels for test samples given a set of labelled training samples as context. Formally, they approximate the posterior predictive distribution:
(5) $p(y_{\text{test}}|\mathbf{x}_{\text{test}},\mathcal{D}_{\text{train}})\approx q_{\theta}(y_{\text{test}}|\mathbf{x}_{\text{test}},\mathcal{D}_{\text{train}})$
where $q_{\theta}$ is parameterised by a transformer-based neural network with parameters $\theta$. For binary empathy classification, the prediction from TabPFN is:
(6) $\hat{y}=\arg\max_{y\in\{0,1\}}q_{\theta}(y|\mathbf{x}_{\text{test}},\mathcal{D}_{\text{train}})$
A simplified representation of the TabPFN architecture is presented in Figure 2 (refer to (Hollmann et al., 2023, 2025) for details). The model includes two separate encoders (X and Y) for processing features and labels, respectively. These encoders also perform basic preprocessing such as data cleaning and input normalisation. The encoded features and labels are concatenated and passed through 12 transformer blocks. Each block includes two multi-head attention mechanisms: one for feature-wise attention and another for item-wise attention. Finally, a decoder transforms the output using a linear layer followed by GELU activation.
3.4. Fine-tuning tabular foundation model
While tabular foundation models provide strong in-context learning performance, their performance can be further enhanced by fine-tuning using the training data. We update the model parameters $\theta$ to $\theta^{\prime}$ by minimising the following binary cross entropy loss with sigmoid activation and temperature scaling:
(7) $\mathcal{L}_{\text{fine}}=-\frac{1}{N_{\text{fine}}}\sum_{i=1}^{N_{\text{fine}}}\left[y_{i}\cdot\log(\sigma(z_{i}^{+}/\tau))+(1-y_{i})\cdot\log(1-\sigma(z_{i}^{+}/\tau))\right]$
where $z_{i}^{+}$ represents the positive class logit for the $i$-th sample, $\sigma(z)=\frac{1}{1+e^{-z}}$ is the sigmoid function, $N_{\text{fine}}$ is the number of samples in fine-tuning, and $\tau$ controls the smoothness of the probability distribution.
To prevent overfitting during fine-tuning, we employ adaptive early stopping and also save the best model based on validation performance.
3.5. Evaluation protocol
We employ two complementary cross-validation approaches: k-fold stratified cross-validation and leave-one-subject-out cross-validation. In each approach, we calculate five evaluation metrics: classification accuracy, precision, recall, F1 score, and area under the curve of the receiver operating characteristic curve (AUC).
K-fold stratified cross-validation is demonstrated as a suitable technique for evaluating models when the number of samples is small. Although the stratified nature maintains the class distribution of the original dataset across all folds, it has a limitation when applied to datasets with multiple samples per subject. Specifically, samples from the same subject may appear in both training and validation sets, potentially leading to an overly optimistic evaluation. This occurs because:
(8) $\begin{split}\exists(\mathbf{x}_{i},y_{i}),&(\mathbf{x}_{j},y_{j})\in\mathcal{D}:\\ \text{subject}(\mathbf{x}_{i})=&\text{subject}(\mathbf{x}_{j})\land(\mathbf{x}_{i},y_{i})\in\mathcal{D}_{\text{train}}^{(f)}\land(\mathbf{x}_{j},y_{j})\in\mathcal{D}_{\text{test}}^{(f)}\end{split}$
To address the aforementioned limitation and evaluate the model’s ability to generalise to unseen individuals, we adopt leave-one-subject-out cross-validation. Let $\mathcal{S}=\{s_{1},s_{2},\ldots,s_{M}\}$ represent the set of $M$ subjects, and $\mathcal{D}_{s_{j}}$ denote the subset of samples from subject $s_{j}$. For each subject $s_{j}$, we define:
(9) $\mathcal{D}_{\text{train}}^{(j)}=\mathcal{D}\setminus\mathcal{D}_{s_{j}},\quad\mathcal{D}_{\text{test}}^{(j)}=\mathcal{D}_{s_{j}}$
This formulation ensures complete separation of subjects between training and test sets:
(10) $\forall(\mathbf{x}_{i},y_{i})\in\mathcal{D}_{\text{train}}^{(j)},(\mathbf{x}_{j},y_{j})\in\mathcal{D}_{\text{test}}^{(j)}:\text{subject}(\mathbf{x}_{i})\neq\text{subject}(\mathbf{x}_{j})$
As stratification cannot be ensured in this cross-validation, some test splits may have all samples from the same class. This is more likely when subjects have a small number of samples. Some evaluation metrics, such as AUROC, are undefined in this scenario. To consider this edge case, We pool the predictions from all subject folds to calculate the evaluation metrics.
4. Experiments
4.1. Dataset
Mathur et al. (2021) released the human-robot interaction (HRI) dataset, which consists of 122 videos (total of 6.9 hours) of interactions between 42 subjects and a robot storyteller. Each interaction followed a controlled protocol where participants listened to three distinct stories, each lasting approximately three minutes. The stories were delivered with randomly assigned narrative voice conditions: half in a first-person narrative voice where the robot told stories about itself, and half in a third-person narrative voice where the robot narrated from an external perspective. After each story, participants completed an 8-item questionnaire using a Likert scale, which produced the ground truth empathy score for each interaction. The labels were binarised – “empathic” and “less-empathic” – using the median empathy score as the threshold. The label distribution is balanced.
For each video, 709 visual behavioural features were extracted using the OpenFace 2.2.0 toolkit (Amos et al., 2016), which captured comprehensive facial and gaze information from human subjects as they listened to the robot’s stories. These features include eye gaze directions, intensity and presence of 17 facial action units, facial landmarks, head pose coordinates and point-distribution model parameters representing facial location, scale, rotation and deformation.
Table 1. Hyperparameter tuning results of 50 trials in Optuna, optimised for 5-fold 10-repeats stratified cross-validated average classification accuracy. Hyperparameters corresponding to the best accuracy are highlighted in bold.
Model Hyperparameter Ranges of values
XGBoost Max depth {3, 4, …, 42, …, 50}
Learning rate [0.01, …, 0.046, …, 0.5]
RF Number of estimators {50, 100, 150, …, 500}
Max depth {3, 4, …, 34, …, 50}
SVM C [0.1, …, 1.0, …, 100] (log scale)
Gamma ‘scale’, [0.001, …, 10] (log scale)
TabPFN Learning rate [1e-6, …, 1e-4, …, 1e-2] (log scale)
Batch size {8, 16, 24, 32, …, 64]
Table 2. 5-fold 10-repeat stratified cross-validation (average $\pm$ standard deviation) scores of tabular foundation models versus baseline models. Single scores refer to the scores reported in the literature. Highest scores are formatted in bold.
Model Accuracy AUC Precision Recall F1 Score
Classical models
XGBoost(Mathur et al., 2021)a 0.69 0.72 0.69 0.69 0.69
XGBoost(Mathur et al., 2021)b $0.598\pm 0.071$ $0.623\pm 0.099$ $0.604\pm 0.079$ $0.589\pm 0.135$ $0.588\pm 0.090$
XGBoostc $0.612\pm 0.093$ $0.651\pm 0.107$ $0.611\pm 0.098$ $0.638\pm 0.161$ $0.615\pm 0.107$
RF $0.625\pm 0.109$ $0.651\pm 0.115$ $0.631\pm 0.119$ $0.639\pm 0.135$ $0.628\pm 0.110$
SVM $0.605\pm 0.080$ $0.641\pm 0.098$ $0.619\pm 0.099$ $0.581\pm 0.119$ $0.592\pm 0.089$
Deep learning
LSTM (Mathur et al., 2021) 0.65 0.71 0.65 0.70
TCN (Mathur et al., 2021) 0.54 0.54 0.76 0.62
TabICL in-context learning $0.634\pm 0.084$ $0.665\pm 0.103$ $0.637\pm 0.100$ $0.660\pm 0.139$ $0.639\pm 0.094$
TabPFN in-context learning $0.622\pm 0.091$ $0.653\pm 0.100$ $0.621\pm 0.096$ $0.651\pm 0.148$ $0.627\pm 0.103$
TabPFN fine-tuned $\textbf{0.715}\pm 0.073$ $0.700\pm 0.091$ $\textbf{0.773}\pm 0.126$ $0.659\pm 0.156$ $\textbf{0.691}\pm 0.096$
  • a
    Scores reported in (Mathur et al., 2021) without specification of metric type (average or best).
  • b
    Our re-implementation based on reported hyperparameters; no code was released for (Mathur et al., 2021).
  • c
    Our implementation with hyperparameter optimisation.
Table 3. Generalisation on unseen subjects through leave-one-subject-out cross-validation. Highest scores are highlighted in bold.
Model Accuracy AUC Precision Recall F1 Score
XGBoost 0.516 0.511 0.519 0.443 0.478
RF 0.533 0.523 0.536 0.492 0.513
SVM 0.590 0.564 0.590 0.590 0.590
TabICL in-context learning 0.541 0.519 0.547 0.475 0.509
TabPFN in-context learning 0.459 0.472 0.451 0.377 0.411
TabPFN fine-tuned (all layers trainable) 0.689 0.642 0.780 0.525 0.627
TabPFN fine-tuned (Frozen X and Y encoders) 0.730 0.669 0.818 0.590 0.686
4.2. Implementation details
We use PyTorch 2.2.0, scikit-learn 1.6.1 and other libraries in Python 3.10.12 to implement our codebase. Details of other libraries and their versions are provided in our GitHub repository. Experiments are run on a single AMD Instinct™ MI250X GPU.
Hyperparameters are tuned using Optuna (Akiba et al., 2019) with its default sampler (tree-structured Parzen estimator) and pruner (median stopping rule). For a fair comparison, we optimise two key hyperparameters for 50 trials for all baseline models and TabPFN fine-tuning setup, with an objective of maximising the average classification accuracy across 5-fold 10-repeats stratified cross-validation. We did not tune the hyperparameters of TabICL and TabPFN in the ICL setup; instead, we used their default configurations. Table 1 presents the list of hyperparameters and corresponding models. The same set of selected hyperparameters is then applied in the leave-one-subject-out experiments to experiment generalisation to unseen subjects.
Tabular foundation models. We use the official pre-trained model release of TabPFN v2222https://huggingface.co/Prior-Labs/TabPFN-v2-clf and TabICL [^3] https://huggingface.co/jingang/TabICL-clf foundation models. Since TabICL’s training code remains unreleased as of 07 April 2025, we use it in the in-context learning experiment only, whereas we experiment with TabPFN in both in-context learning and fine-tuning settings.
Baseline models. We benchmark against XGBoost, which Mathur et al. (2021) reported as the best model for the HRI dataset. Since no official implementation of their XGBoost usage was released, we re-implemented their XGBoost configuration using the three hyperparameters specified in (Mathur et al., 2021). We use default values of other hyperparameters unspecified in (Mathur et al., 2021). We denote this version as XGBoost(Mathur et al., 2021) throughout the paper. Since our implementation of XGBoost(Mathur et al., 2021) failed to reproduce the performance metrics reported in (Mathur et al., 2021), presumably due to undisclosed hyperparameter or training details, we implemented an additional XGBoost model with hyperparameter optimisation via Optuna. Apart from XGBoost, we further implemented random forest (RF) and support vector machine (SVM) algorithms with their default hyperparameters as baseline models.
4.3. Results: Competitive performance on established evaluation protocol
Following the 5-fold stratified 10-repeats cross-validation evaluation protocol established by Mathur et al. (2021), Table 2 compares cross-validated average and standard deviation scores of in-context learning and fine-tuning of tabular foundation models against the baseline models as well as the performance reported in the literature. Since TabPFN’s recommended input capacity is up to 500 features (Hollmann et al., 2025), we select 500 features. For a fair comparison, we also select 500 features for baseline models, including XGBoost, RF and SVM.
Tabular foundation models in ICL setup surpass or achieve comparable performance to baseline models across all evaluation metrics. Assuming the scores reported for XGBoost(Mathur et al., 2021) are average across 50 folds, the fine-tuned TabPFN still outperforms it in classification accuracy, precision and F1 scores. The discrepancy between the reported scores for XGBoost(Mathur et al., 2021) and our re-implementation is likely due to undisclosed hyperparameters or training details.
TabICL outperforms TabPFN. Consistent with the findings of (Qu et al., 2025), our experiment confirms the superior performance of TabICL over TabPFN in the ICL setting. Note that we did not perform any hyperparameter tuning for the ICL experiments.
Fine-tuning tabular foundation model further improves performance. When fine-tuned, TabPFN shows a substantial performance gain and surpasses TabICL. It is worthwhile to note that this improvement comes at the cost of additional computation.
Tabular foundation models outperform other deep learning approaches. As shown in Table 2, tabular foundation models outperform LSTM and TCN deep learning architectures. Their lower performance may be attributed to the use of raw temporal sequences as input, rather than statistical attributes.
Table 4. TabPFN fine-tuning results (5-fold 10-repeats stratified subject-unaware cross-validation) with different experimental settings. Bold indicates the selected configuration.
Configuration Accuracy AUC Precision Recall F1 Score
Optimiser
Schedulerfree (Defazio et al., 2024) $0.704\pm 0.078$ $0.685\pm 0.102$ $0.770\pm 0.122$ $0.615\pm 0.158$ $0.666\pm 0.112$
AdamW $0.699\pm 0.077$ $0.688\pm 0.100$ $0.743\pm 0.121$ $0.650\pm 0.164$ $0.675\pm 0.103$
AdamW + OneCycleLR $0.701\pm 0.077$ $0.689\pm 0.095$ $0.757\pm 0.115$ $0.624\pm 0.148$ $0.669\pm 0.102$
Batch size
8 $0.701\pm 0.077$ $0.689\pm 0.095$ $0.757\pm 0.115$ $0.624\pm 0.148$ $0.669\pm 0.102$
32 $0.708\pm 0.078$ $0.696\pm 0.091$ $0.757\pm 0.116$ $0.656\pm 0.178$ $0.680\pm 0.118$
Frozen layers (Batch size: 8)
None $0.691\pm 0.078$ $0.665\pm 0.108$ $0.743\pm 0.118$ $0.623\pm 0.179$ $0.657\pm 0.120$
X Encoder, Y encoder, Transformer encoder $0.639\pm 0.084$ $0.654\pm 0.102$ $0.675\pm 0.133$ $0.595\pm 0.159$ $0.615\pm 0.104$
X Encoder $0.683\pm 0.079$ $0.661\pm 0.107$ $0.741\pm 0.117$ $0.602\pm 0.190$ $0.640\pm 0.126$
Y encoder $0.692\pm 0.071$ $0.672\pm 0.100$ $0.742\pm 0.111$ $0.620\pm 0.157$ $0.659\pm 0.103$
Transformer encoder $0.643\pm 0.086$ $0.654\pm 0.102$ $0.678\pm 0.132$ $0.599\pm 0.164$ $0.618\pm 0.107$
X Encoder, Y encoder $0.694\pm 0.074$ $0.674\pm 0.099$ $0.751\pm 0.116$ $0.610\pm 0.159$ $0.657\pm 0.104$
Softmax temperature (Batch size: 8)
1.0 $0.691\pm 0.078$ $0.665\pm 0.108$ $0.743\pm 0.118$ $0.623\pm 0.179$ $0.657\pm 0.120$
0.9 $0.683\pm 0.082$ $0.667\pm 0.110$ $0.748\pm 0.123$ $0.591\pm 0.178$ $0.638\pm 0.125$
1.1 $0.690\pm 0.080$ $0.669\pm 0.106$ $0.735\pm 0.123$ $0.623\pm 0.182$ $0.655\pm 0.122$
1.5 $0.683\pm 0.072$ $0.665\pm 0.107$ $0.730\pm 0.116$ $0.615\pm 0.179$ $0.646\pm 0.122$
Number of features selected (Batch size: 8)
25 $0.691\pm 0.078$ $0.665\pm 0.108$ $0.743\pm 0.118$ $0.623\pm 0.179$ $0.657\pm 0.120$
100 $0.680\pm 0.065$ $0.669\pm 0.089$ $0.741\pm 0.108$ $0.593\pm 0.171$ $0.637\pm 0.108$
200 $0.691\pm 0.059$ $0.688\pm 0.090$ $0.731\pm 0.104$ $0.648\pm 0.141$ $0.671\pm 0.076$
500 $0.701\pm 0.077$ $0.689\pm 0.095$ $0.757\pm 0.115$ $0.624\pm 0.148$ $0.669\pm 0.102$
4.4. Results: Improved generalisation on unseen subjects
Given the presence of multiple samples per subject (122 samples from 45 subjects), we adopt a train-test split that accounts for subject identity. Specifically, we use leave-one-subject-out cross-validation, where the splitting is based on participant ID available on the dataset. As this approach introduces no randomness, the results reported in Table 3 are deterministic and presented without any standard deviation.
Generalisation on unseen subjects is more challenging, yet more practical. Models that did well under basic cross-validation – without subject-specific separation – struggled in this setting, which suggests the increased difficulty of the task. Since the model is tested on subjects whose data were entirely absent during training, we avoid inflated estimates of generalisation. The evaluation metrics now better reflect the model’s ability to learn subject-agnostic patterns, which is more realistic for practical deployment. It is also worth noting that while ICL does not involve supervised training, the model still accesses the training data during the forward pass as part of the in-context learning paradigm.
TabPFN ICL yields suboptimal performance. While TabPFN ICL performed competitively under basic cross-validation, it performed poorly in the leave-one-subject-out (LOSO) setting. With a baseline accuracy of 0.5 from a majority-class classifier (i.e., a classifier that predicts all samples as either “empathic” or “less empathic”), TabPFN ICL achieved only 0.459 accuracy, which suggests it made more incorrect than correct predictions. Between TabPFN and TabICL, the latter performs better in the ICL setting, consistent with the results from the subject-unaware cross-validation.
Fine-tuning gives a significant boost to TabPFN cross-subject generalisation. Its classification accuracy improved from 0.459 to 0.730 after fine-tuning. Although the model had no access to the data from the test subject, it effectively learned patterns from other subjects, which suggests good cross-subject generalisation.
4.5. Ablation study
We report a series of experiments on TabPFN fine-tuning (Table 4). Key hyperparameters explored include learning rate, batch size and optimiser choice. We tried the recent scheduler-free optimiser (Defazio et al., 2024) as well as the standard AdamW optimiser, with and without a learning rate scheduler. Since catastrophic forgetting is a challenge when fine-tuning pre-trained models, we tried a warm-up strategy (where the learning rate is linearly increased from a small initial value) with the AdamW optimiser. Specifically, we used a one-cycle learning rate scheduler with cosine annealing. Overall, TabPFN showed minimal performance variations across these optimiser choices, with AdamW and one-cycle scheduler showing the best results. Like optimiser choice, batch size has a modest effect on performance. Increasing it from 8 to 32 improved classification accuracy from 0.701 to 0.708 and AUC from 0.689 to 0.696.
We also tested freezing some pre-trained layers of TabPFN. Freezing the X and Y encoders led to improved performance and faster fine-tuning, so we retained them as frozen and keeping the remaining layers trainable.
We further varied the softmax temperature and the number of input features. Using 500 features led to a slight performance improvement. A softmax temperature of 1.0 (i.e., no temperature scaling) yielded the best classification accuracy (Table 4). Based on these findings, the final model uses the AdamW optimiser with a one-cycle scheduler, a batch size of 32, frozen X and Y encoders, a softmax temperature of 1.0, and 500 selected input features.
Figure 3. In-context learning performance of TabICL and TabPFN across different proportions of in-context data (i.e., training data used in the single forward pass of in-context learning). Results are based on 5-fold, 10-repeat stratified cross-validation without subject group separation.
Figure 3 shows the in-context learning performance of TabICL and TabPFN across varying amounts of context data. In the ICL setup, the context data – i.e., training data – is fed into the model during the single forward pass. Both models generally exhibit improved performance as the amount of context data increases, which suggests the potential for further gains with additional data. However, TabPFN’s AUC peaked when using 80% of the available context data.
4.6. Limitations
We demonstrate the use of tabular foundation models on statistical attributes calculated from temporal video series. While effective, this approach may limit the model’s ability to capture patterns present in the raw temporal data. Incorporating the full temporal sequence into the tabular foundation model could potentially reveal additional patterns and further improve classification performance.
5. Conclusion
This paper investigates the potential of two recent tabular foundation models for detecting empathy elicited through narrative storytelling. Given that raw video sequences are often unavailable due to privacy and ethical constraints, our approach focuses on learning from visual features. We evaluate both in-context learning (ICL) and fine-tuning approaches. Under the published evaluation protocol, both ICL and fine-tuning are competitive and outperform several strong baselines. In addition to the established protocol, we are the first to evaluate cross-subject generalisation in detecting empathy elicited through storytelling. Most models, including the ICL setup of the tabular foundation model, struggle in this setting. However, fine-tuning leads to significantly better generalisation across unseen subjects. Given the continued limitations in accessing raw video data due to privacy and ethical constraints, our results demonstrate practical values and future potential of tabular foundation models.
6. Acknowledgments
This work was supported by resources provided by the Pawsey Supercomputing Research Centre with funding from the Australian Government and the Government of Western Australia.