### v2

# TFMPathy: Tabular Foundation Model for Privacy-Aware, Generalisable Empathy Detection from Videos

Abstract
Detecting empathy from video interactions is an emerging area of research, particularly in healthcare and social robotics. However, privacy and ethical concerns often prevent the release of raw video data, with many datasets instead shared as pre-extracted tabular features. Previous work on such datasets has established classical tree-based models as the state of the art. Motivated by recent successes of large-scale foundation models for text, we investigate the potential of tabular foundation models (TFMs) for empathy detection from video-derived tabular data. Our proposed system, TFMPathy, is demonstrated with two recent TFMs (TabPFN v2 and TabICL) under both in-context learning and fine-tuning paradigms. On a public human-robot interaction benchmark, TFMPathy significantly improves empathy detection accuracy reported in the literature. While the established evaluation protocol in the literature does not ensure cross-subject generalisation, our evaluation scheme also captures such generalisation. We show that TFMPathy under a fine-tuning setup has better cross-subject generalisation capacity over baseline methods (accuracy: $0.590\rightarrow 0.730$; AUC: $0.564\rightarrow 0.669$). Given the ongoing privacy and ethical constraints around raw video sharing, the proposed TFMPathy system provides a practical and scalable path toward building AI systems dependent on human-centred video datasets. Our code is publicly available at https://github.com/hasan-rakibul/TFMPathy (will be made available upon acceptance of this paper).
Index Terms:
Empathy detection, human-robot interaction, tabular data, foundation model, video features, privacy, generalisation.
publicationid: pubid: This work has been submitted to the IEEE for possible publication. Copyright may be transferred without notice.
I Introduction
Figure 1: Overview of the task and motivation. We study the task of detecting empathy of subjects listening to narrative storytelling, where the listener self-reports their empathy level. From the listener’s video, visual features (e.g., eye gaze, facial expressions) are released as a tabular dataset [1]. In this context, we propose TFMPathy, which uses tabular foundation models (TFMs) as the backbone. Our results show that TFMPathy in both in-context learning (ICL) and fine-tuning approaches improves performance, and fine-tuning provides stronger generalisation to unseen subjects.
Empathy, as defined in social psychology [2, 3], refers to an individual’s capacity to perceive, comprehend and vicariously experience the emotional state of others. This interpersonal skill has practical significance in a variety of interaction contexts, such as human-robot interactions (HRIs), clinical communication and educational settings. Training intervention to nurture this skill requires measurement [^1] We use the terms ‘measurement’, ‘detection’ and ‘prediction’ interchangeably. of empathy. Current assessment approaches rely on self-reported questionnaires or labour-intensive manual analysis by external observers [4]. Automated empathy detection through machine learning techniques has gained momentum over the past decade [5]. While the majority of this progress has occurred within natural language processing contexts, i.e., detecting empathy in textual content, empathy detection from video data remains relatively under-explored [5]. Towards this direction, this paper measures empathy from video-based behavioural cues.
Several empathy detection benchmarks are proposed in the literature aiming at different aspects of empathy, such as assessing the level of empathy in an interaction [6, 1], measuring empathic similarity between two contents [7], measuring the changes of someone’s emotional state [8], and detecting the direction of empathy [9]. Our work is grounded on the assessment of the level of empathy, being the most common form in the literature [5]. In particular, we detect the level of empathy elicited by storytelling, i.e., measuring how much empathy a listener demonstrates towards others portrayed in a story [1, 10] (Figure 1).
Being an emerging area of research, the amount of public video datasets is very limited [5]. Datasets of raw video footage featuring people’s activity are difficult to make public due to privacy and ethical concerns. Consequently, the majority of these few existing public datasets, such as human-robot interactions [1, 11], are released as extracted feature vectors (i.e., tabular data) instead of raw video footage. Since privacy and ethical concerns are likely to persist, one of the key challenges in developing an empathy detection system from video is machine learning from tabular datasets.
Traditional approaches to tabular data have long been dominated by tree-based methods like gradient-boosted decision trees [12, 13]. While deep learning has revolutionised learning from texts and images, it has struggled with tabular data due to the heterogeneity between datasets and the heterogeneous nature of the raw data itself. Foundation models like large language models (LLMs) and large vision models (LVMs) have been widely explored in downstream tasks on texts and videos. Similarly, LLMs have also been explored with tabular data [14, 15], but these approaches convert table rows into textual formats such as sentences. As LLMs do not natively support numerical tabular data, a new paradigm of the foundation model has emerged: the tabular foundation model (TFM) [^2] Not to be confused with LLMs. While both are pre-trained foundation models, TFMs are trained with numerical tabular data, whereas LLMs are trained with texts. TFMs are significantly smaller than LLMs. For example, TabPFN has only 7.2 million trainable parameters, in contrast to billions in typical LLMs.. The first one of its kind, a prior data-fitted network (PFN) [16] trained on lots of synthetic tabular data, is TabPFN [12]. Following the success of TabPFN, TabICL (tabular in-context learning) [17] is another TFM, introduced recently in 2025. TFMs natively support in-context learning (ICL), which is a paradigm in which a pre-trained foundation model is conditioned on a small set of input-output examples at inference time. With these examples, the model adapts to new tasks without any gradient updates (i.e., training).
Among the few video-based public empathy detection benchmarks, [1] features empathy elicited by a robot storyteller. In this HRI dataset, human subjects attend to a robot telling different stories that may elicit empathy in the human subjects. Although there are multiple sessions per subject, prior work has failed to consider scenarios where the same subject’s data may appear in both training and test sets. Given that empathy is subjective, such overlaps do not ensure true generalisation performance on unseen subjects. To address this, we design experiments and report results in a new evaluation scheme on this benchmark dataset, where it is ensured that the same subject’s data is not shared between the training and test sets. Our proposed fine-tuning of the TFM achieves impressive results on this cross-subject generalisation (Section IV-D).
Novelty and contributions
Previous work on video-based elicited empathy detection demonstrated classical machine learning models as state-of-the-art approaches. In contrast, our proposed TFMPathy system explores the potential of TFMs and outperforms existing methods by a significant margin (Section IV-C and Section IV-D). We are the first to fine-tune TabPFN to video-based empathy detection with privacy constraints. To make fine-tuning effective and practical, we address key challenges such as catastrophic forgetting and conduct extensive ablation studies covering layer freezing, optimiser and scheduler choices, batch size, softmax temperature and feature dimensionality (Section IV-E). While TFMs are designed for numerical tabular data, we extend them to visual features, in a similar spirit to many works on using LLMs on new domains [18, 19]. We provide new insights and practical strategies for adapting TFMs to privacy-constrained, real-world data derived from videos. On the public HRI dataset, we establish a cross-subject generalisation benchmark, which has not been attempted before. Our key contributions are threefold:
  1. 1.
    We explore the novel application of two TFMs (TabPFN and TabICL) in the ICL setup for empathy detection from video-derived visual features under privacy constraints.
  2. 2.
    Our proposed TFMPathy system introduces a practical and effective fine-tuning strategy for the popular TabPFN foundation model for empathy detection, which demonstrates further performance improvement than ICL, especially in detecting the empathy of unseen subjects.
  3. 3.
    We enhance current evaluation protocols of a public empathy detection benchmark by introducing a subject-independent evaluation benchmark to ensure cross-subject generalisation and to enable fairer and unbiased assessment for real-world applicability.
II Related work
II-A Empathy detection
There are several empathy detection studies in text domains, such as written essays [6, 20], social media conversations [21] and email correspondence [22]. In detecting empathy from videos, [1] released a human-robot interaction (HRI) dataset, aiming to measure empathy elicited by a robot storyteller. They evaluated 10 different machine learning approaches, including eight classical models (e.g., AdaBoost, bagging, decision trees, support vector machine, logistic regression, random forest and XGBoost) and two deep learning architectures (long short-term memory network and temporal convolutional network). For classical machine learning approaches, features were represented as fixed-length vectors of statistical attributes (mean, median, standard deviation and autocorrelation). For deep learning approaches, the features were maintained as temporal sequences. Through 5-fold stratified cross-validation with 10 repeats, XGBoost appeared as the highest-performing classifier. Contrary to their work, we propose ICL and fine-tuning of TFMs in detecting elicited empathy.
Other studies on video-based empathy detection include interaction between human subjects and virtual agents [23, 24], teacher and students [25], human subject and avatar [26], counsellor and clients [27], two human subjects [8], and human and AI agents [28]. In contrast to these empathy modelling approaches, we focus on measuring elicited empathy from the public HRI benchmark.
II-B Application of foundation models
Textual foundation models (i.e., LLMs) have been explored in several stages of empathy detection from textual data. For example, [29] used GPT-4 and Llama-3 LLMs, and [30] used GPT series LLMs in a zero-shot setting to predict empathy in written essays. On the same task, [30] and [31] explored few-shot prompt engineering (i.e., ICL) with GPT LLMs. Apart from the prediction, LLMs are also explored in data augmentation [32, 33] and data annotation [29, 31].
Motivated by the success of LLMs in empathy detection from text, we explore two recent TFMs released in 2025 – TabPFN v2 [13] and TabICL [17] – for tabular video features. While TFMs’ ICL capacity has been demonstrated in numerical datasets such as OpenML test suites [34], we explore its capability in dealing with video features. In addition to ICL, we fine-tune TabPFN v2, which further boosts empathy classification performance.
III Method
Figure 2: Overview of our proposed TFMPathy pipeline for empathy detection. Temporal multimodal features (e.g., facial landmarks, eye gaze) are first extracted for each sample $\mathbf{X}_{i}\in\mathbb{R}^{T_{i}\times d}$. These sequences are summarised by computing $n$ statistical attributes (e.g., mean, median) over time, resulting in a fixed-size tabular representation $\mathbf{X}_{i}\in\mathbb{R}^{n\times d}$. This representation is used for binary empathy classification $y_{i}\in\{0,1\}$) via either (top) in-context learning, where a frozen tabular foundation model (TFM) predicts labels with a single forward pass over the training data, or (bottom) fine-tuning of the foundation model with frozen X and Y encoders. The model predicts whether a subject is empathic or less empathic.
Figure 2 presents an overview of our approach for empathy detection using TFMs. Temporal multimodal features, such as facial landmarks and eye gaze, are first aggregated into fixed-size tabular representations via summary statistics. These are then used for binary empathy classification through either ICL or fine-tuning of a TFM. Further details of each component are provided in the following subsections.
III-A Problem formulation
We formulate empathy detection in HRI videos either as a supervised learning or an ICL problem. Given a dataset $\mathcal{D}=\{(\mathbf{X}_{i},y_{i})\}_{i=1}^{N}$, consisting of $N$ interaction samples from $M$ subjects, where $\mathbf{X}_{i}\in\mathbb{R}^{T_{i}\times d}$ represents a $d$-dimensional feature vector across $T_{i}$ frames of the $i$-th video sample and $y_{i}\in\{0,1\}$ denotes the corresponding binary empathy label.
III-B Data preprocessing for tabular foundation model
The feature vector $\mathbf{X}_{i}$ contains frame-level visual behavioural cues (e.g., eye gaze directions, facial action unit intensities, facial landmarks, and other non-verbal indicators) extracted using toolkits such as OpenFace [35]. These temporal features can be processed in two ways: (1) time-series modelling or (2) transformed into fixed-length representations by computing statistical attributes (e.g., mean, median, standard deviation and autocorrelation) across the temporal axis. Prior work [1] demonstrated that the latter approach works better for empathy detection in the HRI context.
From $\mathbf{X}_{i}\in\mathbb{R}^{T_{i}\times d}$, we first filter out temporal points having zeros in all features, and then calculate statistical attributes. For $n$ statistical attributes, the input feature representation becomes a fixed-size representation $\mathcal{X}_{i}\in\mathbb{R}^{nd}$. For reference, in OpenFace features (709 dimensions) and four statistical attributes, the resultant feature becomes a vector of $709\times 4=2836$ elements. From this, we first drop dimensions where it is constant across all samples.
Both TabPFN and TabICL support up to 500 features, and therefore, we further reduce the feature dimension. We rank all features by their ANOVA F‑values with the target labels on the training set and retain the top $d^{\prime}$ features. At evaluation time, we apply this fixed subset to each test split, without recomputing F‑values to prevent any information leakage from test to training. This feature selection process is performed independently within each fold of cross-validation.
With this tabular representation, our goal is to learn a function $f:\mathbb{R}^{d^{\prime}}\rightarrow[0,1]$ that maps the input features to a conditional probability distribution over the target variable (empathy level). For a general supervised learning task, it becomes:
$p(y=1|\mathbf{x};\theta)=f_{\theta}(\mathbf{x})$ (1)
where $\mathbf{x}\in\mathcal{X}$ represents the input features, and $\theta$ denotes the model parameters optimised during training.
III-C Traditional supervised learning approaches
In supervised learning, we use labelled training data $\mathcal{D}_{\text{train}}$ to optimise model parameters:
$\theta^{*}=\arg\min_{\theta}\mathcal{L}(\theta,\mathcal{D}_{\text{train}})$ (2)
where $\mathcal{L}$ is a loss function (e.g., binary cross-entropy) and $\theta^{*}$ represents the optimised parameters. During inference, predictions are made on a separate test set:
$p(y=1|\mathbf{x};\theta^{*})=f_{\theta^{*}}(\mathbf{x})$ (3)
III-D In-context learning with tabular foundation model
In-context learning explicitly utilises the training data in the inference framework:
$p(y=1|\mathbf{x},\mathcal{D}_{\text{train}};\theta)=f_{\theta}(\mathbf{x},\mathcal{D}_{\text{train}})$ (4)
TFMs such as TabPFN and TabICL, by design, support ICL. These foundation models are pre-trained on millions of synthetic tabular datasets to predict labels for test samples, given a set of labelled training samples as context. Formally, they approximate the posterior predictive distribution:
$p(y_{\text{test}}|\mathbf{x}_{\text{test}},\mathcal{D}_{\text{train}})\approx q_{\theta}(y_{\text{test}}|\mathbf{x}_{\text{test}},\mathcal{D}_{\text{train}})$ (5)
where $q_{\theta}$ is parameterised by a transformer-based neural network with parameters $\theta$. For binary empathy classification, the prediction from TabPFN is:
$\hat{y}=\arg\max_{y\in\{0,1\}}q_{\theta}(y|\mathbf{x}_{\text{test}},\mathcal{D}_{\text{train}})$ (6)
A simplified representation of the TabPFN architecture is included in Figure 2 (refer to [12, 13] for details). The model includes two separate encoders (X and Y) for processing features and labels, respectively. These encoders also perform basic preprocessing such as data cleaning and input normalisation. The encoded features and labels are concatenated and passed through 12 transformer blocks. Each block includes two multi-head attention mechanisms: one for feature-wise attention and another for item-wise attention. Finally, a decoder transforms the output using a linear layer followed by GELU activation.
III-E Fine-tuning tabular foundation model
While TFMs provide strong ICL performance, their performance can be further enhanced by fine-tuning using the training data. We update the model parameters $\theta$ to $\theta^{\prime}$ by minimising the following binary cross-entropy loss with sigmoid activation and temperature scaling:
$\mathcal{L}_{\text{fine}}=-\frac{1}{N_{\text{fine}}}\sum_{i=1}^{N_{\text{fine}}}\Bigl{[}y_{i}\cdot\log(\sigma(z_{i}^{+}/\tau))\\ +(1-y_{i})\cdot\log(1-\sigma(z_{i}^{+}/\tau))\Bigr{]}$ (7)
where $z_{i}^{+}$ represents the positive class logit for the $i$-th sample, $\sigma(z)=\frac{1}{1+e^{-z}}$ is the sigmoid function, $N_{\text{fine}}$ is the number of samples in fine-tuning, and $\tau$ controls the smoothness of the probability distribution.
Catastrophic forgetting is a common challenge with fine-tuning pre-trained models. To address this, we apply a warm-up strategy (where the learning rate is linearly increased from a small initial value) with the optimiser. Specifically, we used a one-cycle learning rate scheduler with cosine annealing. To prevent overfitting during fine-tuning, we employ adaptive early stopping, which dynamically adjusts the patience parameter (i.e., the number of steps before stopping the training) throughout training rather than relying on a fixed value. We save the best model based on validation performance.
III-F Evaluation protocol
We employ two complementary cross-validation approaches: k-fold stratified cross-validation and leave-one-subject-out cross-validation. In each approach, we calculate five evaluation metrics: classification accuracy, precision, recall, F1 score, and area under the curve of the receiver operating characteristic curve (AUC).
K-fold stratified cross-validation is demonstrated as a suitable technique for evaluating models when the number of samples is small. Although the stratified nature maintains the class distribution of the original dataset across all folds, it has a limitation when applied to datasets with multiple samples per subject. Specifically, samples from the same subject may appear in both training and validation sets, potentially leading to an overly optimistic evaluation. This occurs because:
$\begin{split}\exists(\mathbf{x}_{i},y_{i}),&(\mathbf{x}_{j},y_{j})\in\mathcal{D}:\\ \text{subject}(\mathbf{x}_{i})=&\text{subject}(\mathbf{x}_{j})\land(\mathbf{x}_{i},y_{i})\in\mathcal{D}_{\text{train}}^{(f)}\land(\mathbf{x}_{j},y_{j})\in\mathcal{D}_{\text{test}}^{(f)}\end{split}$ (8)
To address the aforementioned limitation and evaluate the model’s ability to generalise to unseen individuals, we adopt leave-one-subject-out cross-validation. Let $\mathcal{S}=\{s_{1},s_{2},\ldots,s_{M}\}$ represent the set of $M$ subjects, and $\mathcal{D}_{s_{j}}$ denote the subset of samples from subject $s_{j}$. For each subject $s_{j}$, we define:
$\mathcal{D}_{\text{train}}^{(j)}=\mathcal{D}\setminus\mathcal{D}_{s_{j}},\quad\mathcal{D}_{\text{test}}^{(j)}=\mathcal{D}_{s_{j}}$ (9)
This formulation ensures complete separation of subjects between training and test sets:
$\forall(\mathbf{x}_{i},y_{i})\in\mathcal{D}_{\text{train}}^{(j)},(\mathbf{x}_{j},y_{j})\in\mathcal{D}_{\text{test}}^{(j)}:\text{subject}(\mathbf{x}_{i})\neq\text{subject}(\mathbf{x}_{j})$ (10)
As stratification cannot be ensured in this cross-validation, some test splits may have all samples from the same class. This is more likely when subjects have a small number of samples. Some evaluation metrics, such as AUROC, are undefined in this scenario. To consider this edge case, we pool the predictions from all subject folds to calculate the evaluation metrics.
IV Experiments and Results
IV-A Dataset
[1] released the human-robot interaction (HRI) dataset, which consists of 122 videos (a total of 6.9 hours) of interactions between 42 subjects and a robot storyteller. Each interaction followed a controlled protocol where participants listened to three distinct stories, each lasting approximately three minutes. The stories were delivered with randomly assigned narrative voice conditions: half in a first-person narrative voice where the robot told stories about itself, and half in a third-person narrative voice where the robot narrated from an external perspective. After each story, participants completed an 8-item questionnaire using a Likert scale, which produced the ground truth empathy score for each interaction. [1] binarised the labels into “empathic” and “less-empathic” categories using the median empathy score as a threshold. The label distribution is balanced.
For each video, 709 visual behavioural features were extracted using the OpenFace 2.2.0 toolkit [35], which captured comprehensive facial and gaze information from human subjects as they listened to the robot’s stories. These features include eye gaze directions, intensity and presence of 17 facial action units, facial landmarks, head pose coordinates and point-distribution model parameters representing facial location, scale, rotation and deformation.
TABLE I: Hyperparameter tuning results of 50 trials in Optuna, optimised for 5-fold 10-repeats stratified cross-validated average classification accuracy. Hyperparameters corresponding to the best accuracy are highlighted in bold.
Model Hyperparameter Ranges of values
XGBoost Max depth {3, 4, …, 42, …, 50}
Learning rate [0.01, …, 0.046, …, 0.5]
RF Number of estimators {50, 100, 150, …, 500}
Max depth {3, 4, …, 34, …, 50}
SVM C [0.1, …, 1.0, …, 100] (log scale)
Gamma ‘scale’, [0.001, …, 10] (log scale)
TabPFN Learning rate [1e-6, …, 1e-4, …, 1e-2] (log scale)
Batch size {8, 16, 24, 32, …, 64]
TABLE II: 5-fold 10-repeat stratified cross-validation (average $\pm$ standard deviation) scores of our proposed TFMPathy versus baseline models. Single scores refer to the scores reported in the literature. Highest scores are formatted in bold. ICL refers to in-context learning, and FT refers to fine-tuning.
Model Accuracy AUC Precision Recall F1 Score
Classical models
XGBoost[1]a 0.69 0.72 0.69 0.69 0.69
XGBoost[1]b $0.598\pm 0.071$ $0.623\pm 0.099$ $0.604\pm 0.079$ $0.589\pm 0.135$ $0.588\pm 0.090$
XGBoostc $0.612\pm 0.093$ $0.651\pm 0.107$ $0.611\pm 0.098$ $0.638\pm 0.161$ $0.615\pm 0.107$
RF $0.625\pm 0.109$ $0.651\pm 0.115$ $0.631\pm 0.119$ $0.639\pm 0.135$ $0.628\pm 0.110$
SVM $0.605\pm 0.080$ $0.641\pm 0.098$ $0.619\pm 0.099$ $0.581\pm 0.119$ $0.592\pm 0.089$
Deep learning
LSTM [1] 0.65 0.71 0.65 0.70
TCN [1] 0.54 0.54 0.76 0.62
TFMPathyICL(TabICL) $0.634\pm 0.084$ $0.665\pm 0.103$ $37\pm 0.100 0.660$$60\pm 0.139$ $0.639\pm 0.094$
TFMPathyICL(TabPFN) $0.622\pm 0.091$ $0.653\pm 0.100$ $0.621\pm 0.096$ $0.651\pm 0.148$ $0.627\pm 0.103$
TFMPathyFT(TabPFN) $\textbf{0.715}\pm 0.073$ $0.700\pm 0.091$ $\textbf{0.773}\pm 0.126$ $0.659\pm 0.156$ $\textbf{0.691}\pm 0.096$
  • a
    Scores reported in [1] without specification of metric type (average or best).
  • b
    Our re-implementation based on reported hyperparameters; no code was released for [1].
  • c
    Our implementation with hyperparameter optimisation.
IV-B Implementation details
We use PyTorch 2.2.0, scikit-learn 1.6.1 and other libraries in Python 3.10.12 for experiments. Details of other libraries and their versions are provided in our public GitHub repository. Experiments are run on a single AMD Instinct™ MI250X GPU (64 GB GCD).
Hyperparameters are tuned using Optuna [36] with its default sampler (tree-structured Parzen estimator) and pruner (median stopping rule). For a fair comparison, we optimise two key hyperparameters for 50 trials for all baseline models and TabPFN fine-tuning setup, with an objective of maximising the average classification accuracy across 5-fold 10-repeats stratified cross-validation. We did not tune the hyperparameters of TabICL and TabPFN in the ICL setup; instead, we used their default configurations. Table I presents the list of hyperparameters and corresponding models. The same set of selected hyperparameters is then applied in the leave-one-subject-out experiments to experiment with generalisation to unseen subjects.
IV-B1 Tabular foundation models
We use the official pre-trained release of TabPFN v2333https://huggingface.co/Prior-Labs/TabPFN-v2-clf and TabICL [^4] https://huggingface.co/jingang/TabICL-clf foundation models. Since TabICL’s training code was unavailable, we use it in the ICL experiment only, whereas we experiment with TabPFN in both ICL and fine-tuning settings.
IV-B2 Baseline models
We benchmark against XGBoost, which [1] reported as the best model for the HRI dataset. Since no official implementation of their XGBoost usage was released, we re-implemented their XGBoost configuration using the three hyperparameters specified in [1]. We use the default values of other hyperparameters unspecified in [1]. We denote this version as XGBoost[1] throughout the paper. Since our implementation of XGBoost[1] failed to reproduce the performance metrics reported in [1], presumably due to undisclosed hyperparameters or training details, we implemented an additional XGBoost model with hyperparameter optimisation via Optuna. Apart from XGBoost, we further implemented random forest (RF) and support vector machine (SVM) algorithms with their default hyperparameters as baseline models.
The work proposing TabPFN benchmarks against gradient boosting methods, including XGBoost, across 18 numerical datasets, because gradient boosting methods continue to dominate tabular classification tasks [12]. We propose a privacy-aware framework to address the unavailability of raw videos due to privacy and ethics constraints. In such settings, tabular representations of pre-extracted visual features are often the only available modality. Prior research [1] on the HRI benchmark also established XGBoost as the best method, outperforming deep learning models. While advanced video models in this dataset could capture more expressive temporal cues, that would require access to full video sequences, which is unavailable to “protect privacy” [1]. Therefore, the above selection of baseline models is appropriate, given the tabular nature of the benchmark.
IV-C Results: Competitive performance on established evaluation protocol
Following the 5-fold stratified 10-repeats cross-validation evaluation protocol established by [1], Table II compares cross-validated average and standard deviation scores of ICL and fine-tuning of TFMPathy against the baseline models as well as the performance reported in the literature. Since TabPFN’s recommended input capacity is up to 500 features [13], we select 500 features. For a fair comparison, we also select 500 features for baseline models, including XGBoost, RF and SVM.
TFMPathy in ICL setup surpasses or achieves comparable performance to baseline models across all evaluation metrics. Assuming the scores reported for XGBoost[1] are average across 50 folds, the fine-tuned TabPFN still outperforms it in classification accuracy, precision and F1 scores. The discrepancy between the reported scores for XGBoost[1] and our re-implementation is likely due to undisclosed hyperparameters or training details.
TabICL outperforms TabPFN. Consistent with the findings of [17], our experiment confirms the superior performance of TabICL over TabPFN in the ICL setting. Note that we did not perform any hyperparameter tuning for the ICL experiments.
Fine-tuning tabular foundation model further improves performance. When fine-tuned, TabPFN shows a substantial performance gain and surpasses TabICL. This improvement, however, comes at a cost of additional computation. TabPFN has a total of 7.2 million trainable parameters, and in our experiment on a 64 GB GPU, the average fine‑tuning time per step ranged from 3.86 to 4.08 seconds for each of the 50 folds.
TFMPathy outperform other deep learning approaches. As shown in Table II, TFMPathy outperform LSTM and TCN deep learning architectures. Their lower performance may be attributed to the use of raw temporal sequences as input, rather than statistical attributes.
TABLE III: Generalisation on unseen subjects through leave-one-subject-out cross-validation. Highest scores are highlighted in bold. ICL refers to in-context learning, and FT refers to fine-tuning.
Model Accuracy AUC Precision Recall F1 Score
XGBoost 0.516 0.511 0.519 0.443 0.478
RF 0.533 0.523 0.536 0.492 0.513
SVM 0.590 0.564 0.590 0.590 0.590
TFMPathyICL(TabICL) 0.541 0.519 0.547 0.475 0.509
TFMPathyICL(TabPFN) 0.459 0.472 0.451 0.377 0.411
TFMPathyFT(TabPFN; all layers trainable) 0.689 0.642 0.780 0.525 0.627
TFMPathyFT(TabPFN; frozen X and Y encoders) 0.730 0.669 0.818 0.590 0.686
IV-D Results: Improved generalisation on unseen subjects
Given the presence of multiple samples per subject (122 samples from 45 subjects), we adopt a train-test split that accounts for subject identity. Specifically, we use leave-one-subject-out cross-validation, where the splitting is based on participant ID available on the dataset. As this approach introduces no randomness, the results reported in Table III are deterministic and presented without any standard deviation.
Generalisation on unseen subjects is more challenging, yet more practical. Models that did well under basic cross-validation – without subject-specific separation – struggled in this setting, which suggests the increased difficulty of the task. Since the model is tested on subjects whose data were entirely absent during training, we avoid inflated estimates of generalisation. The evaluation metrics now better reflect the model’s ability to learn subject-agnostic patterns, which is more realistic for practical deployment. It is also worth noting that while ICL does not involve supervised training, the model still accesses the training data during the forward pass as part of the in-context learning paradigm.
TabPFN ICL yields suboptimal performance. While TabPFN ICL performed competitively under basic cross-validation, it performed poorly in the leave-one-subject-out (LOSO) setting. With a baseline accuracy of 0.5 from a majority-class classifier (i.e., a classifier that predicts all samples as either “empathic” or “less empathic”), TabPFN ICL achieved only 0.459 accuracy, which suggests it made more incorrect than correct predictions. Between TabPFN and TabICL, the latter performs better in the ICL setting, consistent with the results from the subject-unaware cross-validation.
Fine-tuning gives a significant boost to TabPFN cross-subject generalisation. Its classification accuracy improved from 0.459 to 0.730 after fine-tuning. Although the model had no access to the data from the test subject, it effectively learned patterns from other subjects, which suggests good cross-subject generalisation. For each of the 42 subjects, the average fine‑tuning time per step ranged from 4.70 to 4.98 seconds.
The primary architectural difference between TabPFN and TabICL lies in their attention mechanisms: TabICL uses a column-then-row attention scheme, which makes it faster than TabPFN. Since TabICL outperformed TabPFN in the ICL setup, it is reasonable to expect that fine-tuning TabICL could further improve performance, which we leave for future work.
TABLE IV: Fine-tuning of TFMPathy(TabPFN) results (5-fold 10-repeats stratified subject-unaware cross-validation) with different experimental settings. Bold indicates the selected configuration.
Configuration Accuracy AUC Precision Recall F1 Score
Optimiser
Schedulerfree [37] $0.704\pm 0.078$ $0.685\pm 0.102$ $0.770\pm 0.122$ $0.615\pm 0.158$ $0.666\pm 0.112$
AdamW $0.699\pm 0.077$ $0.688\pm 0.100$ $0.743\pm 0.121$ $0.650\pm 0.164$ $0.675\pm 0.103$
AdamW + OneCycleLR $0.701\pm 0.077$ $0.689\pm 0.095$ $0.757\pm 0.115$ $0.624\pm 0.148$ $0.669\pm 0.102$
Batch size
8 $0.701\pm 0.077$ $0.689\pm 0.095$ $0.757\pm 0.115$ $0.624\pm 0.148$ $0.669\pm 0.102$
32 $0.708\pm 0.078$ $0.696\pm 0.091$ $0.757\pm 0.116$ $0.656\pm 0.178$ $0.680\pm 0.118$
Frozen layers (Batch size: 8)
None $0.691\pm 0.078$ $0.665\pm 0.108$ $0.743\pm 0.118$ $0.623\pm 0.179$ $0.657\pm 0.120$
X encoder, Y encoder, Transformer encoder $0.639\pm 0.084$ $0.654\pm 0.102$ $0.675\pm 0.133$ $0.595\pm 0.159$ $0.615\pm 0.104$
X encoder $0.683\pm 0.079$ $0.661\pm 0.107$ $0.741\pm 0.117$ $0.602\pm 0.190$ $0.640\pm 0.126$
Y encoder $0.692\pm 0.071$ $0.672\pm 0.100$ $0.742\pm 0.111$ $0.620\pm 0.157$ $0.659\pm 0.103$
Transformer encoder $0.643\pm 0.086$ $0.654\pm 0.102$ $0.678\pm 0.132$ $0.599\pm 0.164$ $0.618\pm 0.107$
X encoder, Y encoder $0.694\pm 0.074$ $0.674\pm 0.099$ $0.751\pm 0.116$ $0.610\pm 0.159$ $0.657\pm 0.104$
Softmax temperature (Batch size: 8)
1.0 $0.691\pm 0.078$ $0.665\pm 0.108$ $0.743\pm 0.118$ $0.623\pm 0.179$ $0.657\pm 0.120$
0.9 $0.683\pm 0.082$ $0.667\pm 0.110$ $0.748\pm 0.123$ $0.591\pm 0.178$ $0.638\pm 0.125$
1.1 $0.690\pm 0.080$ $0.669\pm 0.106$ $0.735\pm 0.123$ $0.623\pm 0.182$ $0.655\pm 0.122$
1.5 $0.683\pm 0.072$ $0.665\pm 0.107$ $0.730\pm 0.116$ $0.615\pm 0.179$ $0.646\pm 0.122$
Number of features selected (Batch size: 8)
25 $0.691\pm 0.078$ $0.665\pm 0.108$ $0.743\pm 0.118$ $0.623\pm 0.179$ $0.657\pm 0.120$
100 $0.680\pm 0.065$ $0.669\pm 0.089$ $0.741\pm 0.108$ $0.593\pm 0.171$ $0.637\pm 0.108$
200 $0.691\pm 0.059$ $0.688\pm 0.090$ $0.731\pm 0.104$ $0.648\pm 0.141$ $0.671\pm 0.076$
500 $0.701\pm 0.077$ $0.689\pm 0.095$ $0.757\pm 0.115$ $0.624\pm 0.148$ $0.669\pm 0.102$
IV-E Ablation study
Table IV reports a series of experiments on fine-tuning TabPFN (TFMPathyFT). Key hyperparameters explored include learning rate, batch size and optimiser choice. We tried the recent scheduler-free optimiser [37] as well as the standard AdamW optimiser, with and without a learning rate scheduler. Overall, TabPFN showed minimal performance variations across optimiser choices, with AdamW and one-cycle scheduler showing the best results. Like optimiser choice, batch size has a modest effect on performance. Increasing it from 8 to 32 improved classification accuracy from 0.701 to 0.708 and AUC from 0.689 to 0.696.
We also tested freezing three pre-trained layers of TabPFN: X encoder of 768 trainable parameters, Y encoder of 576 parameters, and Transformer encoder of 7.1 million trainable parameters. Freezing the X and Y encoders led to improved performance, likely because they are already well-generalised from TabPFN’s large-scale pre-training. Freezing preserves generalisable representations and enables faster fine-tuning, so we kept the X and Y encoders frozen and kept the remaining layers trainable.
We further varied the softmax temperature and the number of input features. Using 500 features led to a slight performance improvement. A softmax temperature of 1.0 (i.e., no temperature scaling) yielded the best classification accuracy (Table IV). Based on these findings, the final model uses the AdamW optimiser with a one-cycle scheduler, a batch size of 32, frozen X and Y encoders, a softmax temperature of 1.0, and 500 selected input features.
Figure 3: In-context learning performance of TFMPathy across different proportions of in-context data (i.e., training data used in the single forward pass of in-context learning). Results are based on 5-fold, 10-repeat stratified cross-validation without subject group separation.
Figure 3 shows the ICL performance of the TFMPathy system using both TabICL and TabPFN across varying amounts of context data. In the ICL setup, the context data – i.e., training data – is fed into the model during the single forward pass. Both models generally exhibit improved performance as the amount of context data increases, which suggests the potential for further gains with additional data. However, TabPFN’s AUC peaked when using 80% of the available context data.
IV-F Limitations
We demonstrate the use of TFMs on statistical attributes calculated from temporal video series. While effective, this approach may limit the model’s ability to capture patterns present in the raw temporal data. Incorporating the full temporal sequence into the TFM could potentially reveal additional patterns and further improve classification performance.
V Conclusion
This paper proposes TFMPathy, a system utilising the potential of TFMs for detecting empathy elicited through narrative storytelling. Given that raw video sequences are often unavailable due to privacy and ethical constraints, TFMPathy learns from visual features and does not require any raw videos. Experimented on two recent TFMs (TabICL and TabPFN), we evaluate both ICL and fine-tuning approaches. In a public evaluation benchmark on human-robot interaction, TFMPathy in both ICL and fine-tuning settings is competitive and outperforms several strong baselines. In addition to the established protocol, we are the first to evaluate cross-subject generalisation in detecting empathy elicited through storytelling. While most models, including TFMPathy in the initial ICL setup, struggle in generalisation, TFMPathy in the fine-tuning setup leads to significantly better generalisation across unseen subjects. Given the persistent privacy and ethical constraints in sharing raw videos, our proposed TFMPathy system offers practical value and future potential for AI systems that rely on human-centred video datasets.
Acknowledgments
This work was supported by resources provided by the Pawsey Supercomputing Research Centre with funding from the Australian Government and the Government of Western Australia.